OpenStudy (anonymous):

Can someone plz help i will medal and fan explain the steps in solving the following absolute value equation. 2[x-2]=12

OpenStudy (anonymous):

divide by two:(x-2)=6 do the absolute value: x+2=6 x=4

OpenStudy (anonymous):

tward can i ask you another question?

OpenStudy (anonymous):

tag me in another question and I will gladly

OpenStudy (anonymous):

how doi tag lol

OpenStudy (anonymous):

@Lord_Frieza

OpenStudy (welshfella):

there is another solution to the equation the value of x - 2 can also be equal to -6

OpenStudy (anonymous):

TWard3212 How do you show where an inequality is true on a number line

OpenStudy (welshfella):

so we have x - 2 =6 or x - 2 = -6 so x = 8 , -4

OpenStudy (welshfella):

x = 4 is incorrect

OpenStudy (anonymous):

thank you